OpenStudy (anonymous):

Find the values of x on the interval [−π, π] where the tangent line to the graph of y = sin(x) cos(x) is horizontal. (Enter your answers as a comma-separated list.)

OpenStudy (rogue):

A trig identity is \[a \sin u \cos u = \frac {a}{2} \sin (2u)\]So you can write your equation as\[y = \sin(x) \cos(x) = \frac {1}{2} \sin (2x)\]Use the crain rule here\[y' = \frac {d}{dx} \frac {1}{2}\sin (2x) = \frac {1}{2} \cos (2x) \frac {d}{dx} 2x = \cos (2x)\]The curve will have horizontal tangents when y' = 0.\[y' = 0 = \cos (2x)\]On the interval [-pi, pi], solution to that is\[x = \pm \frac {\pi}{4}, \pm \frac {3 \pi}{4}\]

OpenStudy (anonymous):

there is too many lines in your answer i cant nderstand it

OpenStudy (rogue):

Too many lines... :/ What part don't you get? Also, by 'crain rule' I meant chain rule.

OpenStudy (anonymous):

what is the answer, it looks like this frac//sin{}[2xs[frac {dx) d and a whole lot of weird lines

OpenStudy (rogue):

:o oh, I think it isn't loading properly for you. Try refreshing the page or giving it more time. Answer is: x = - pi/4, + pi/4, - 3pi/4, +3pi/4

OpenStudy (rogue):

Here is a picture of what I typed.

OpenStudy (anonymous):

ahhh gotcha, well you got it right, good job :)